OpenStudy (anonymous):

how to solve the integration of y*(9-y^2)^1/2

OpenStudy (ujjwal):

put y^2=t

OpenStudy (anonymous):

there is a sqrt on (9-y^2) do i have to solve this question with integeration by parts??

OpenStudy (ujjwal):

not really. if you put y^2=t you have ydy=dt/2 and hence you will have to integrate ((9-t)^1/2)/2 which looks easier to solve

OpenStudy (anonymous):

this was the question of multiple integrals, i have solved one integral and i get this, the anx is 9 bt i am not able to solve it

OpenStudy (ujjwal):

the limits on the right integral are functions of y. That means the limits shows the range of x. therefore integrate at first with dx and then with dy. you should get 9 as an answer finally.

OpenStudy (anonymous):

if i put y^2=t i will also have to change the limits

OpenStudy (anonymous):

yes i integrated it first with dx

OpenStudy (ujjwal):

o to (9-y^2)^1/2 is the range of x

OpenStudy (ujjwal):

one you are done integrating with x, you can put y^2=t and that will not change limit. coz limits are constants.

OpenStudy (ujjwal):

*once

ganeshie8 (ganeshie8):

you may try "9-y^2 = t" also

OpenStudy (ujjwal):

yeah, that is more simpler.

OpenStudy (anonymous):

ok i'll try

OpenStudy (anonymous):

i am not getting the correct anx by putting y^2=t and the same limits

OpenStudy (ujjwal):

what are you getting?

OpenStudy (anonymous):

the anx is in cube root or if i simplify it it is in points -0.087....

OpenStudy (ujjwal):

after putting 9-y^2 =t find dt at first and tell me what you get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-cuberoot 9/3

OpenStudy (anonymous):

just cuberoot 9 /3

OpenStudy (ujjwal):

find dt in terms of dy at first

OpenStudy (anonymous):

dt/dy=-2y i have done the whole method and the limits are not constant, because it is iterated integration

OpenStudy (ujjwal):

the limits after first integration is constant. put 9-y^2=t so, now you need to integrate (sqrt t)/-2 in terms of dt

OpenStudy (ujjwal):

put 9-y^2 =t after you are done integrating with x

OpenStudy (ujjwal):

\[\int_{0}^{3}\int_{0}^{\sqrt{9-y^2}}ydxdy\] \[=\int_{0}^{3}y\sqrt{9-y^2}dy\] Now put \(9-y^2=t\). Therefore \[ydy=-\frac{dt}{2}\] So, now you need to integrate \[\int_{0}^{3}\frac{\sqrt{t}}{-2}dt\] You can integrate this, right?

OpenStudy (anonymous):

yes i did the same the anx is not 9

OpenStudy (ujjwal):

oh, sorry, the limits change. when you put 9-y^2=t, For y=0, you have t=9 and when y=3 then t=0 So, the limit is 9 to 0 you need to integrate \[\int_{9}^{0}\frac{\sqrt{t}}{-2}dt\] you should get 9 now.

OpenStudy (ujjwal):

or else you can switch back to x later on after integration, i.e express t in terms of x again and keep the limits unchanged.